OpenStudy (laddiusmaximus):

rates of change area of a square with a respect to its side s when s=46

OpenStudy (anonymous):

a= s*s da\ds = 2s => 92

OpenStudy (anonymous):

\[A=s^2\] \[A'=2ss'\]

OpenStudy (anonymous):

cannot go further unless you know s' or A'

OpenStudy (anonymous):

it said with respect to side

OpenStudy (anonymous):

so d/ds(s^2) = 2s(ds\ds)